# Enabling Natural Zero-Shot Prompting on Encoder Models via Statement-Tuning

### v2

@@ Appendix B Finetuning Setup
#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze Yahoo Answers Topic Emotion
Llama-2-7b-chat 7B 91.0 67.9 42.8 95.2 82.1 61.9 54.3
Mistral-7B-Instruct-v0.2 7B 93.8 78.2 44.8 96.2 87.0 65.0 57.0
Qwen1.5-7B-Chat 7B 91.4 79.4 43.8 95.1 82.4 63.9 58.0
Pythia-6.7B 6.7B 84.6 66.9 39.2 91.6 74.0 38.3 52.0
Pythia-2.7B 2.7B 80.8 63.5 41.5 90.8 71.7 35.5 47.5
Phi-2 2.7B 90.8 74.0 44.7 93.8 81.6 58.4 58.4
FlanT5-Large 770M 66.2 78.7 39.7 75.3 59.9 38.0 34.6
Qwen1.5-0.5B-Chat 500M 73.4 56.1 38.5 84.2 68.8 36.1 31.4
BART-large-mnli 406M 52.2 32.4 42.0 50.6 51.1 7.1 10.0
FlanT5-Small 60M 52.0 32.6 41.4 75.8 50.0 9.1 9.8
@@
Our Approach: RoBERTa-base (4k) 125M 69.6 69.8 59.3 92.7 82.8 41.2 55.5
Our Approach: RoBERTa-large (Best) 355M 85.0 72.5 74.7 95.3 93.0 51.1 55.1
Our Approach: RoBERTa-large (50k) 355M 85.0 71.4 72.2 95.3 92.9 51.1 55.1
@@
RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2
RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2
Table 5: Comparison of our approach against many pretrained open-source Encoder-Decoder and Decoder-only Pretrained Large Language Models on 7 Natural Language Understanding tasks in 32-shot conditions. We highlight all scores in red where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
@@ Appendix E Task Categories Breakdown
For the statement tuning task diversity, we group corpus based on task categories as follows:
@@
    Sentiment Analysis (SA): Yelp Polarity

### v3

@@ Appendix B Finetuning Setup
#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze Yahoo Answers Topic Emotion Avg
Meta-Llama-3-70B-Instruct 70B 95.2 78.9 46.1 96.6 86.5 66.4 58.8 75.5
Llama-2-13b-chat-hf 13B 93.2 71.6 44.3 95.7 84.9 61.1 56.6 72.5
Llama-2-7b-chat 7B 91.0 67.9 42.8 95.2 82.1 61.9 54.3 70.7
Mistral-7B-Instruct-v0.2 7B 93.8 78.2 44.8 96.2 87.0 65.0 57.0 74.6
Qwen1.5-7B-Chat 7B 91.4 79.4 43.8 95.1 82.4 63.9 58.0 73.4
Pythia-6.7B 6.7B 84.6 66.9 39.2 91.6 74.0 38.3 52.0 63.8
Pythia-2.7B 2.7B 80.8 63.5 41.5 90.8 71.7 35.5 47.5 61.6
Phi-2 2.7B 90.8 74.0 44.7 93.8 81.6 58.4 58.4 71.7
FlanT5-Large 770M 66.2 78.7 39.7 75.3 59.9 38.0 34.6 56.1
Qwen1.5-0.5B-Chat 500M 73.4 56.1 38.5 84.2 68.8 36.1 31.4 55.5
BART-large-mnli 406M 52.2 32.4 42.0 50.6 51.1 7.1 10.0 35.0
FlanT5-Small 60M 52.0 32.6 41.4 75.8 50.0 9.1 9.8 38.7
@@
Our Approach: RoBERTa-base (4k) 125M 75.0(0.6) 70.0(1.9) 60.3(1.0) 92.4(0.8) 79.7(3.5) 38.2(2.5) 45.4(3.2) 65.9
Our Approach: RoBERTa-large (Best) 355M 85.1(0.8) 71.5(1.9) 74.7(1.8) 95.3(0.8) 91.9(0.2) 50.2(2.2) 49.8(1.4) 74.1
Our Approach: RoBERTa-large (10k) 355M 85.1(0.8) 71.5(0.8) 72.6(1.8) 95.3(0.3) 91.0(1.0) 48.2(0.8) 48.4(3.6) 73.1
@@
RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2 -
RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2 -
Table 5: Comparison of our approach against many pretrained open-source encoder-Decoder and Decoder-only Pretrained Large Language Models on 7 Natural Language Understanding tasks in 32-shot conditions. We highlight all scores in gray where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.
@@ Appendix E Task Categories Breakdown
For the statement tuning task diversity, we group datasets based on task categories as follows (evaluation datasets are underlined):
@@
    Sentiment Analysis (SA): Yelp Polarity, Amazon Polarity
@@
  10. 10.
    Sentence Completion: BCOPA, StoryCloze
  11. 11.
    Emotion Recognition: Emotion
  12. 12.
    Topic Classification: Yahoo Answer Topic
  13. 13.
    Nonliteral Reasoning: FigQA
Appendix F N-shot Correlation
Figure 6 shows the correlation of accuracies achieved in different N-shot settings with various shot numbers.
Appendix G Inference Speed Comparison
We report the average examples/sec processed for each of the datasets in Table 6. It is important to note that all models are run on a single GPU, except for Meta-Llama-3-70B-Instruct and Llama-2-13B-chat which were run on 4 and 2 GPUs, respectively.
Model BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion Avg
Qwen1.5-0.5B-Chat 0.1 0.1 0.1 0.1 0.1 0.1 0.1 0.1
phi-2 1.4 1.4 1.4 1.4 1.5 1.4 1.4 1.4
Meta-Llama-3-70B-Instruct* 2.9 1.3 3.2 0.9 4.9 1.1 2.1 2.3
flan-t5-large 8.2 13.2 13.2 13.2 13.2 13.2 13.2 12.5
Llama-2-13b-chat-hf* 8.7 5.7 12.8 4.3 15.7 4.4 6.9 8.3
Our Approach (roberta-large) 9.3 14.5 15.0 15.0 14.7 3.1 5.1 11.0
bart-large-mnli 9.7 14.1 14.0 14.2 14.1 13.7 13.8 13.4
pythia-6.9b 12.0 0.6 4.6 0.4 0.6 2.2 0.4 3.0
Llama-2-7b-chat-hf 12.5 0.6 4.6 0.4 0.6 2.3 0.5 3.1
Mistral-7B-Instruct-v0.2 12.8 0.5 2.7 0.3 0.5 1.7 0.4 2.7
pythia-2.8b 13.6 16.7 24.9 15.2 27.2 15.1 20.9 19.1
flan-t5-small 13.9 39.2 39.1 39.3 39.4 39.3 39.3 35.6
Our Approach (roberta-base) 17.9 49.8 50.0 49.8 49.9 10.3 17.0 34.9
Table 6: The average examples per second processed by each model on each task. * indicates that the model required the use of more than one GPU.
Llama-2-13b Qwen1.5-7B Pythia-6.9B Pythia-2.9B Phi-2 Qwen1.5-0.5B
#Parameters 13B 7B 6.9B 2.9B 2.7B 500M
Quantization 8bit 4bit 4bit 4bit 4bit 4bit
Sequence Length 4096 2048 2048 2048 2048 2048
lora_r 32 32 32 32 32 32
lora_alpha 16 64 64 64 64 64
lora_dropout 0.05 0.05 0.05 0.05 0.05 0.05
adam_beta2 0.999 0.999 0.95 0.95 0.95 0.999
adam_epsilon 1e-8 1e-8 0.00001 0.00001 0.00001 1e-8
max_grad_norm none none 1.0 1.0 1.0 none
optimizer adamw_bnb_8bit adamw_torch adamw_torch adamw_torch adamw_torch adamw_torch
gradient acc. 4 4 4 4 4 4
micro batch size 2 1 1 1 1 1
lr_scheduler cosine cosine cosine cosine cosine cosine
learning_rate 0.0002 0.0002 0.00001 0.00001 0.000003 0.0002
Table 7: Instruction Tuning Hyperparameters
Appendix H Statement-Tuning Data Instruction-Tuned Decoders
As seen in Table 8, we report the comparison of our approach with several decoders that were instruct-tuned using an instruction-tuning dataset created using the same training corpora used for Statement-Tuning. Training details are outlined in Table 7. Furthermore, the templates used to form instructions are based on those used for Flan Wei et al. (2022a) (we make the dataset available here: Anonymized Repository).
The models tested are a subset of the ones reported in Table 1 due to the time and computational expense of instruction-tuning and hardware limitations. However, even when the LLM models are instruction-tuned on the same data as the Statement-Tuned RoBERTa models, we observe similar trends where performance of the RoBERTa-base model tends to match the performance of all models up to 6.9B parameters on all tasks except for BCOPA. Furthermore, the RoBERTa-large model approaches or exceeds performance on all tasks for models with 7B+ parameters. The same trend of dominating performance on FigQA and StoryCloze is observed.
#Parameters BCOPA MRPC FigQA Amazon Polarity StoryCloze YA Topic Emotion Avg
Llama-2-13b 13B 89.6 60.8 40.9 93.7 82.4 53.2 51.6 67.5
Qwen1.5-7B 7B 87.2 78.9 41.4 94.8 75.7 47.8 56.5 68.9
Pythia-6.9B 6.9B 82.8 68.1 40.0 71.7 71.5 16.4 27.5 54.0
Pythia-2.9B 2.9B 79.6 67.9 40.3 77.2 69.7 21.2 30.6 55.2
Phi-2 2.7B 87.2 68.1 41.7 85.6 77.8 38.4 53.5 64.6
Qwen1.5-0.5B 500M 72.4 68.4 39.4 49.8 67.6 33.2 72.4 57.6
[…]
Our Approach: RoBERTa-base (4k) 125M 72.4(0.5) 69.6(1.1) 60.7(0.9) 92.3(0.8) 78.5(2.7) 37.9(2.7) 46.6(4.3) 65.4
Our Approach: RoBERTa-large (Best) 355M 85.1(0.7) 71.8(0.8) 74.2(1.4) 95.4(0.4) 92.1(0.7) 49.9(2.1) 50.7(1.4) 75.3
Our Approach: RoBERTa-large (10k) 355M 85.1(0.7) 71.5(0.8) 73.0(2.4) 95.4(0.4) 91.1(0.8) 48.4(0.7) 49.1(3.2) 73.4
Full/3000-shot:
RoBERTa-base (FT) 125M 74.2 87.0 88.1 94.3 - 71.0 82.2 -
RoBERTa-large (FT) 355M 86.0 87.6 92.0 96.5 - 68.5 78.2 -
Table 8: Comparison of our approach against many pretrained open-source encoder-Decoder and Decoder-only Instruction-tuned Pretrained Large Language Models on 7 Natural Language Understanding tasks in Zero-shot conditions. FT stands for Full Finetuning and is included for reference. For Statement-Tuning, we report the average across 5 training runs and 5 evaluation runs and include the average standard deviation in parenthesis. We highlight all scores in gray where our approach with RoBERTa-base (best) exceeds or is equal to the score given by the model.